A three phase induction motor is wound for 4 poles and is supplied from 50 Hz system. Calculate the rotor frequency when the rotor runs at 600 r.p.m. 


1
Expert's answer
2021-08-31T01:43:00-0400

"N_s=\dfrac{120f}{p}=\dfrac{120\times{50}}{4}=1500\ rpm\\
N_r=N_s(1-s)\\
600=1500(1-s)\\
s=1-\dfrac{6}{15}=0.6\\
Rotor\ frequency\ is\ =sf=0.6\times{50}=30\ Hertz"